OpenStudy (anonymous):

Find a polynomial of degree 2 with integer coefficients and zeros 5-i and 5+i.

OpenStudy (anonymous):

you have lots of choices of methods here

OpenStudy (anonymous):

one way is to multiply out \[(x-(5+i))(x-(5-i))\] which is not as hard as it seems. but is kind of a pain

OpenStudy (anonymous):

next choice is to work backwards you know \[x=5+i\] \[x-5=i\] \[(x-5)^2=i^2=-1\] \[x^2-10x+25=-1\] \[x^2-10x+26=0\] is the equation you started with

OpenStudy (anonymous):

the other is to memorize that if \(a+bi\) is the zero of a polynomial of degree two with leading coefficient 1, that quadratic equation is \[x^2-2ax+(a^2+b^2)=0\]

OpenStudy (anonymous):

pick whichever method you like

OpenStudy (anonymous):

Thank you for the very helpful and speedy response!

OpenStudy (anonymous):

yw